# LADM: Long-context Training Data Selection with Attention-based Dependency Measurement for LLMs

### v1

@@ 6.3 Ablation Study on LADM
To validate the effectiveness of the contextual dependency measurement in LADM, we conduct additional ablation studies on the standard deviation weights in Eq. 4 and the length weights in both Eq. 4 and Eq. 5. Experimental results of Llama2-7B on the LongBench dataset are shown in Table 6. Higher standard deviation weights indicate more diverse dependency patterns within long-context samples and higher length weights focus more on span pairs with greater distances, making them crucial for long-range contextual dependency measurement. Therefore, we observe a significant performance drop without these weights.
@@ 6.4 Observations and Findings
In Figure 4, we show the median scores of ProLong and LADM framework on different categories. The results show that articles (PhilPapers, ArXiv and PubMed Central), legal documents (FreeLaw) and books (PG-19) tend to receive higher scores in our method. This is likely due to their coherent logic and strong interrelations between paragraphs, indicating complex and varied dependencies within these data types.
Moreover, we analyze the features of data with lower scores. Wikipedia data samples often manifest as concatenated paragraphs that may only share a theme, but exhibit weak connectivity between paragraphs. Similarly, the Ubuntu IRC dataset is composed of group chat discussions. These conversations show duplicated patterns and different sections involve different users, time points and topics, resulting in weak contextual coherence. We show samples from different categories in Appendix D.
Compared to our method, ProLong tends to give much higher scores to samples from the Ubuntu IRC dataset containing segments with similar patterns despite low relevance between them. Moreover, samples from the PG-19 dataset get the lowest score, which indicates that ProLong may struggle to capture the deeper contextual relationships between seemingly unrelated sections within books. These sections may be connected through narrative development, which is not apparent through analysis of isolated segments. This highlights the importance of measuring the dependencies within full context rather than just assessing this relationship between individual segments.
@@ Appendix D Data samples
Random 24.89 34.04 62.63 68.36 21.18 42.22
ProLong 26.11 35.84 63.68 68.46 20.56 42.93
LADM 25.33 36.26 64.44 68.64 20.69 43.07
@@
Random 37.83 44.97 74.41 75.15 26.32 51.74
ProLong 39.63 44.80 74.83 75.95 24.97 52.04
LADM 40.96 45.14 74.07 76.06 26.19 52.49
@@
Random 38.17 43.09 69.40 74.65 24.72 50.01
ProLong 44.37 49.40 77.23 78.05 25.09 54.83
LADM 52.76 48.12 75.34 78.51 25.46 56.04
@@
Random 50.53 50.60 76.77 79.36 25.46 56.54
ProLong 50.61 50.00 77.31 79.44 25.95 56.66
LADM 50.37 49.66 77.57 79.92 26.44 56.79

### v2

@@ 6.3 Ablation Study on LADM
To validate the effectiveness of the contextual dependency measurement in LADM, we conduct additional ablation studies on the standard deviation weights $\sigma_{j}$ in Eq. 4 and the length weights $\frac{j-i}{N}$ and $\frac{j}{N}$ in Eq. 4 and Eq. 5. Experimental results of Llama2-7B on the LongBench dataset are shown in Table 6. Higher standard deviation weights indicate more diverse dependency patterns within long-context samples and higher length weights focus more on span pairs with greater distances, making them crucial for long-range contextual dependency measurement. Therefore, we observe a significant performance drop without these weights.
@@ 6.4 Observations and Findings
In Figure 4, we show the median scores of ProLong and LADM framework on different categories. The results show that articles (PhilPapers, ArXiv, and PubMed Central), legal documents (FreeLaw), and books (PG-19) tend to receive higher scores in our method. This is likely due to their coherent logic and strong interrelations between paragraphs, indicating complex and varied dependencies within these data types.
Moreover, we analyze the features of data with lower scores. Wikipedia data samples often manifest as concatenated paragraphs that may only share a theme but exhibit weak connectivity between paragraphs. Similarly, the Ubuntu IRC dataset is composed of group chat discussions. These conversations show duplicated patterns and different sections involve different users, time points and topics, resulting in weak contextual coherence. We show samples from different categories in Appendix D.
Compared to our method, ProLong tends to give much higher scores to samples from the Ubuntu IRC dataset containing segments with similar patterns, despite low relevance between them. Moreover, samples from the PG-19 dataset get the lowest score, which indicates that ProLong may struggle to capture the deeper contextual relationships between seemingly unrelated sections within books. These sections may be connected through narrative development, which is not apparent through analysis of isolated segments. This highlights the importance of measuring the dependencies within the full context rather than just assessing the relationship between individual segments.
@@ 7 Conclusion
Acknowledgments
We thank our colleagues Chen Wang, Chong Li, Zixuan Ren, Yaochen Zhu, and Pu Jian for their insightful and constructive feedback. Furthermore, we thank all reviewers for their detailed reviews and valuable comments. This work is supported by the National Key R&D Program of China No.2022ZD0160602 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grant No.XDA04080400. This work is also supported by Zhongguancun Academy Project No.20240103.
@@ Appendix D Data samples
Random 24.89 34.04 62.63 68.36 21.18 42.22 (-1.14)
ProLong 26.11 35.84 63.68 68.46 20.56 42.93 (-0.43)
LADM 25.33 36.26 64.44 68.64 20.69 43.07 (-0.29)
@@
Random 37.83 44.97 74.41 75.15 26.32 51.74 (-0.81)
ProLong 39.63 44.80 74.83 75.95 24.97 52.04 (-0.51)
LADM 40.96 45.14 74.07 76.06 26.19 52.49 (-0.06)
@@
Random 38.17 43.09 69.40 74.65 24.72 50.01 (-10.44)
ProLong 44.37 49.40 77.23 78.05 25.09 54.83 (-5.62)
LADM 52.76 48.12 75.34 78.51 25.46 56.04 (-4.41)
@@
Random 50.53 50.60 76.77 79.36 25.46 56.54 (-0.22)
ProLong 50.61 50.00 77.31 79.44 25.95 56.66 (-0.10)
LADM 50.37 49.66 77.57 79.92 26.44 56.79 (+0.03)